ligoatindacut:

what is the value of x in the equation 2/3(1/2x 12)=1/2(1/3x 14)-3

Tranquility:

Hello! Welcome to QuestionCove! Please post your complete question to receive better assistance. Your question isn't complete as an operation is missing in the parenthesis. Is it (1/2x - 12)? Or (1/2x + 12)? Regardless, the first step to solving your question is to use the distributive property a(b + c) = ab + ac a(b - c) = ab - ac Distribute the number outside of the parenthesis on both sides. After that, you can add/subtract so that way you get your 'x' on one side of the equal sign and a number on the other side and you can solve for 'x' by multiplying/dividing to finally isolate 'x'